abowl contain 10 chips, of which 8 are marked $ 2 each and 2 are marke $5 each, let aperson choose at random and without replcement , 3 chips from this bowl, if the person is to receive the sum of the resulting amount
find his expectation

Answer by stanbon(48568)

abowl contain 10 chips, of which 8 are marked $ 2 each and 2 are marke $5 each, let aperson choose at random and without replcement , 3 chips from this bowl, if the person is to receive the sum of the resulting amount
find his expectation
--------------------
There are 10C3 = 120 groups of 3 chips.
-----------
8C3=56 of them have $2 marked on all three; value = $6 with probability of (56/120)
---------------
[8C2][2C1]=28*2=56 of them have two chips varked with $2 and one chip marked
with a $5; value of each group is $9 with probability of (56/120)
--------------
[8C1][2C2]=8 of them have one marked $2 and two marked $5; value of each
group is $12 withprobability of (8/120)
----------
Expected value is $6*(56/120)+$9*9+(8/120)*$12=$7.80
Cheers,
Stan H.